# Evaluating Pre-trained Convolutional Neural Networks and Foundation Models as Feature Extractors for Content-based Medical Image Retrieval

### v1

@@ Abstract
Medical image retrieval refers to the task of finding similar images for given query images in a database, with applications such as diagnosis support, treatment planning, and educational tools for inexperienced medical practitioners. While traditional medical image retrieval was performed using clinical metadata, content-based medical image retrieval (CBMIR) relies on the characteristic features of the images, such as color, texture, shape, and spatial features. Many approaches have been proposed for CBMIR, and among them, using pre-trained convolutional neural networks (CNNs) is a widely utilized approach. However, considering the recent advances in the development of foundation models for various computer vision tasks, their application for CBMIR can be also investigated for its potentially superior performance.
In this study, we used several pre-trained feature extractors from well-known pre-trained CNNs (VGG19, ResNet-50, DenseNet121, and EfficientNetV2M) and pre-trained foundation models (MedCLIP, BioMedCLIP, OpenCLIP, CONCH and UNI) and investigated the CBMIR performance on a subset of the MedMNIST V2 dataset, including eight types of 2D and 3D medical images. Furthermore, we also investigated the effect of image size on the CBMIR performance.
Our results show that, overall, for the 2D datasets, foundation models deliver superior performance by a large margin compared to CNNs, with UNI providing the best overall performance across all datasets and image sizes. For 3D datasets, CNNs and foundation models deliver more competitive performance, with CONCH achieving the best overall performance. Moreover, our findings confirm that while using larger image sizes (especially for 2D datasets) yields slightly better performance, competitive CBMIR performance can still be achieved even with smaller image sizes. Our codes to generate and reproduce the results are available at: https://github.com/masih4/MedImageRetrieval.
@@ 1 Introduction
Content-based medical image retrieval (CBMIR) refers to the application of content-based retrieval in the medical imaging domain. CBMIR can assist medical experts in making image-based diagnoses. For example, in the context of histological images, pathologists might distinguish between cancerous and non-cancerous whole slide images by comparing the image under analysis with the top-$k$ relevant images returned by the CBMIR model ($k$ can be defined by the user, e.g., 3 or 5). Besides this application, CBMIR can be used for other purposes, such as image-based training of medical professionals or practitioners and treatment planning [3].
@@
A standard CBIR or CBMIR model typically consists of three main building blocks including an image pre-processing pipeline, feature extraction, and feature similarity measurements between query image features and stored image features in the database using various distance-based methods. While early models primarily relied on image-processing-based features such as color, texture, shape, or spatial features, most recent models use automatically extracted features from pre-trained models, such as pre-trained convolutional neural networks (CNNs) or vision transformers (ViTs). Many pre-trained CNNs, such as ResNet family models, EfficientNet family models, or ViT family models, have been exploited as retrieval feature extractors in previous studies [4, 5, 6]. In general, compared to CBIR, which has been widely studied, less research has been conducted on CBMIR, and most recent CBMIR studies are based on CNNs as feature extractors [7, 8].
@@
Besides standard pre-trained CNNs and ViTs, which are typically trained on the ImageNet dataset [9] in a supervised manner, there has been a significant amount of research focused on foundation models recently. A foundation model refers to a large-scale pre-trained model that can be used for various applications, typically in computer vision or natural language processing. Foundation models are usually trained in a self-supervised or unsupervised manner using approaches such as contrastive language-image pre-training (CLIP) [10] or knowledge distillation with no Labels (DINO) [11, 12]. Various foundation models have been developed or applied for tasks in medical computer vision, and some very recent studies have used a selection of them for radiological image retrieval [6, 7, 13]. Another understudied area in medical image retrieval, especially for foundation models, is the investigation of the effect of image size on retrieval performance. Most foundation models were originally trained on 224$\times$224 images, and the effect of using smaller-sized images on their general performance has not been well-investigated.
@@
In this study, we conducted a comprehensive analysis of CBMIR for various types of 2D and 3D radiological and non-radiological medical images, using both pre-trained CNNs (as benchmarks) and recently developed foundation models that have not been thoroughly investigated in previous studies. We evaluated their retrieval performance for zero-shot retrieval without any post-processing, fine-tuning, or retraining. Furthermore, we investigated the effect of image size (ranging from very small images with only 28 pixels in one dimension to standard 224-pixel images) on retrieval performance across eight types of 2D and 3D medical images.
@@ 2.1 Datasets
We used a subset of eight datasets introduced in the MedMNIST V2 study [14, 15]. We selected datasets from various imaging modalities, including colon pathology, dermatoscopy, chest X-ray, fundus imaging, breast ultrasound, blood cell microscopy, abdominal CT, and electron microscopy, to investigate image retrieval performance across a wide range of medical images, both in 2D (six datasets) and 3D (two datasets). The basic statistics of these datasets, including the imaging type (2D or 3D), number of training/testing images per dataset, number of classes, and imaging modality are listed in Table 1, and example images from each dataset are shown in Figure 1. In the latest version of MedMNIST V2, 2D datasets are provided in various sizes of 28$\times$28, 64$\times$64, 128$\times$128, and 224$\times$224 pixels, and 3D datasets are provided in two sizes of 28$\times$28$\times$28 and 64$\times$64$\times$64 pixels. Further details about each sub-dataset can be found in the respective publications as referenced in Table 1. It should be noted that we evaluated the performance of the retrieval model on the test images, and the training images were used to build the database for image search.
@@ 2.2 Pre-trained models as feature extractors
We used both well-known pre-trained CNN models (as benchmark) and a number of foundation models as feature extractors to compare their retrieval performance on the selected sub-datasets from the MedMNIST V2. The pre-trained CNN models include VGG19 [22], ResNet50 [23], DenseNet121 [24], and EfficientNetV2M [25], while the pre-trained foundation models comprise MedCLIP [26], BioMedCLIP [27], OpenCLIP [28, 29], CONCH [30], and UNI [31]. For foundation model selection, we chose models that were not thoroughly investigated in previous studies for various types of medical image retrieval. We excluded some foundation models such as DINO V1 [11], DINO V2 [12], SAM [32], MEDSAM [33], Virchow [34], and MAE [35], as well as pre-trained ViT-based models like the original ViT [36] or Swin Transformers [37], because they have already shown inferior performance for medical image retrieval compared to the foundation models selected for this study [7, 31, 6]. The basic statistics of the selected CNN and foundation models are listed in Table 2, and a brief description of each is provided in the following.
Table 2: Basic statistics of the utilized CNN (upper part) and foundation models (lower part) in this study. Nat. = Natural, Med. = Medical.
@@ 2.2.4 EfficientNetV2M
EfficientNet is a family of CNN models initially designed to systematically and uniformly scale all dimensions of the model (depth, width, and resolution) using compound coefficients [43]. In the original study, neural architecture search [44] was employed to find a baseline model (EfficientNetB0) consisting of convolutional and MB convolutional blocks [45], which was then scaled up to larger models for better performance (EfficientNetB1-B7). In the second version of the EfficientNet family (EfficientNetV2), a combination of Fused-MB convolutional blocks [46] (in the initial stages) and MB convolutional blocks (in the later stages) was employed and the network was optimized for improving training speed, parameter efficiency and performance. Other techniques, such as non-uniform scaling strategy, progressive learning (gradually increasing the image resolution during training), and exploiting regularizers (dropout, RandAugment, mixup), were also utilized to further boost the performance. In this study, we utilized EfficientNetV2M, which has been shown to deliver excellent performance for e.g. image classification and retrieval [25, 47] while being computationally efficient. Like other CNN-based models, we used average pooling to convert the feature maps from 8$\times$8$\times$1280 to 1280 features.
@@ 2.2.5 MedCLIP
Vision-language models such as CLIP [10] have shown excellent performance for a variety of computer vision tasks. CLIP consists of two main building blocks, an image encoder and a text encoder that map the input images and texts to a common latent space. Using contrastive learning, the model is trained to maximize the similarity between paired images and texts. The original CLIP model was trained on 400 million general image-text pairs (named as WIT-400M) extracted from the internet. However, in the context of medical data, finding large image-text paired datasets is very challenging. In contrast to CLIP, MedCLIP is trained with three types of datasets, including medical image-text, only medical image, and only medical text datasets. To train the model, two similarity matrices were created: one referred to as the semantic similarity matrix and the other as the predicted similarity matrix. To build the semantic similarity matrix, the MetaMap [48] approach was used to extract entities from texts and images in all three types of datasets. To build the predicted similarity matrix, a text encoder (BioClinicalBERT [^1] https://huggingface.co/emilyalsentzer/Bio_ClinicalBERT) and an image encoder (Swin Transformer [37]) were used. The utilized image encoder was pre-trained on the ImageNet dataset. Finally, to train the model, semantic matching loss was used instead of InfoNCE contrastive loss used in the CLIP study. The model was pre-trained using two X-ray datasets, namely, MIMIC-CXR (around 600K image-text pairs) and CheXpert (around 223K images). To extract features using the pre-trained MedCLIP model, we used the vision encoder embedding, which yields 512 features for each input image.
@@ 2.2.6 BioMedCLIP
This model is an extension of the CLIP study, specifically designed for medical data. BioMedCLIP pre-training was performed using the PMC-15M dataset, which contains around 15 million image-text pairs automatically extracted from 4.4 million PubMed articles. BioMedCLIP employs a vision transformer (ViT-B/16-224) [36] as the image encoder and PubMedBERT [49] with an extended token size of 256 as the text encoder. To extract image features from BioMedCLIP, we used the image encoder, which yields 512 features for each input image.
@@ 2.2.8 CONCH
Contrastive learning from captions for histopathology or CONCH is a vision-language foundation model trained on over 1.17 million histological image-text pairs sourced from in-house educational notes and the PubMed central open access dataset. Similar to pre-trained CNNs, which are trained on natural images and used for various applications such as medical image analysis, we also considered using CONCH and UNI (see next section) as feature extractors for a variety of medical image types. The dataset creation to train the CONCH model involved three main steps. First, using a YOLO V5 model [52], histological images were detected. Then a GPT-style model was utilized to split the image captions [53], and finally, a CLIP model was employed to match the images with their corresponding captions. The model training was conducted using the CoCa approach [54], which includes an image encoder (ViT-B/16), a GPT-style text encoder, and a GPT-style text decoder. Unlike CLIP, which only uses contrastive loss to maximize cosine similarity scores for pairing images and text, CoCa is also equipped with a captioning objective to generate proper text for input images. The vision encoder of the model was pre-trained on 16 million in-house images using the iBOT self-supervised approach [55], while the language model was pre-trained using more than 550,000 surgical histopathology reports from Massachusetts General Hospital and over 400,000 PubMed abstracts related to histopathology. For our study, we used the vision encoder of the CONCH model to generate 512 features for the input images.
@@ 2.2.9 UNI
UNI is a large-scale visual pre-trained model, originally trained on histological images. UNI was trained using the Mass-100K dataset, which contains more than 100 million H&E-stained image patches extracted from over 100,000 whole slide images. The network training was based on the DINO V2 [12] model, a self-supervised, student-teacher-based approach for pre-training. For a given image, we extracted features using the ViT-G/14 encoder of the UNI model, which yielded a feature vector of size 1024.
@@ 2.3 Distance measurement
As mentioned in Section 2.2, we used pre-trained CNNs and foundation models for feature extraction from both 2D and 3D datasets and all these models were originally trained on 2D images. For 2D datasets, we simply used the 2D training and testing images for feature extraction. For 3D datasets, however, we extracted features from each slice of a 3D train/test volume and concatenated these features to obtain the feature map for the entire 3D volume. While this approach can be computationally expensive, we chose to use it because our initial results indicated that applying feature reduction techniques such as principal component analysis (PCA) [56], autoencoders [57], t-Distributed Stochastic Neighbor Embedding (t-SNE) [58], or uniform manifold approximation and projection (UMAP) [59] led to a decrease in performance (refer to Table S5 and Table S6 in the supplementary materials). To measure the distance between the extracted features of the query images/volumes and the stored features in the database, we used the well-known cosine similarity index, as suggested by previous studies on medical image retrieval [6, 10]. After calculating the index for all images (query samples and all samples in the training set), we ranked them, and the top-$k$ most similar images were selected as the output of the retrieval model.

### v2

@@ Abstract
Medical image retrieval refers to the task of finding similar images for given query images in a database, with applications such as diagnosis support. While traditional medical image retrieval relied on clinical metadata, content-based medical image retrieval (CBMIR) depends on image features, which can be extracted automatically or semi-automatically. Many approaches have been proposed for CBMIR, and among them, using pre-trained convolutional neural networks (CNNs) is a widely utilized approach. However, considering the recent advances in the development of foundation models for various computer vision tasks, their application for CBMIR can also be investigated.
In this study, we used several pre-trained feature extractors from well-known pre-trained CNNs and pre-trained foundation models and investigated the CBMIR performance on eight types of two-dimensional (2D) and three-dimensional (3D) medical images. Furthermore, we investigated the effect of image size on the CBMIR performance.
Our results show that, overall, for the 2D datasets, foundation models deliver superior performance by a large margin compared to CNNs, with the general-purpose self-supervised model for computational pathology (UNI) providing the best overall performance across all datasets and image sizes. For 3D datasets, CNNs and foundation models deliver more competitive performance, with contrastive learning from captions for histopathology model (CONCH) achieving the best overall performance. Moreover, our findings confirm that while using larger image sizes (especially for 2D datasets) yields slightly better performance, competitive CBMIR performance can still be achieved even with smaller image sizes. Our codes to reproduce the results are available at: https://github.com/masih4/MedImageRetrieval.
@@ 1 Introduction
Content-based medical image retrieval (CBMIR) refers to the application of CBIR in the medical imaging domain. Compared to other medical computer vision tasks such as classification or segmentation, less research has been conducted on CBMIR. However, CBMIR models have a distinct advantage over classification models, for example. The output of a classification model for a given query image consists of probabilities for each class and, usually, through thresholding, provides a determined label for the test image. On the other hand, the output of a CBMIR model for a given query image consists of the top-$k$ relevant images ($k$ can be defined by the user, e.g., 3 or 5), usually retrieved from a large database of images. In many cases, the database also stores clinical metadata or diagnostic reports alongside the images, which can complement the model output as auxiliary information [3]. As a result, CBMIR provides medical experts and biologists with more comprehensive insights into diverse patterns and features, which allows them to identify the most similar patterns across different images. In contrast to classification models, which only produce labels for given images as black-box models, CBMIR keeps clinicians and medical experts in the loop for grading or classification, which is often necessary for artificial intelligence applications in healthcare. CBMIR models facilitate the diagnostic workflow by providing additional information; however, medical experts are still required to provide the final diagnostic conclusion rather than replacing medical experts with a machine learning or deep learning model [4, 5]. Besides these applications, CBMIR can also be used for other purposes, such as image-based training of medical professionals and treatment planning [6]. From a technical perspective, one advantage of CBMIR over classification or segmentation is that it is usually performed in an unsupervised manner, whereas classification and segmentation models typically require labeled data for training or fine-tuning.
@@
A standard CBIR or CBMIR model typically consists of three main building blocks: an image pre-processing pipeline, feature extraction, and feature similarity measurements between query images and stored images in the database using various distance-based methods. While early models primarily relied on handcrafted features including color, texture, shape, and spatial information, most recent models use automatically extracted features from pre-trained models, such as pre-trained convolutional neural networks (CNNs) or vision transformers (ViTs), which are referred to as deep features. Although deep features suffer from low explainability and interpretability, they have demonstrated superior performance in many computer vision tasks, including image retrieval, due to their better generalization capability [7, 8, 9]. Many pre-trained CNNs, such as ResNet family models, EfficientNet family models, or ViT family models, have been exploited as retrieval feature extractors in previous studies [10, 11, 12]. In general, compared to CBIR, which has been widely studied, less research has been conducted on CBMIR, and the majority of recent CBMIR studies rely on CNNs as feature extractors [13, 14].
@@
Besides standard pre-trained CNNs and ViTs, which are typically trained on the ImageNet dataset [15] in a supervised manner, there has been a significant amount of research focused on foundation models recently, which can also extract deep features. A foundation model refers to a large-scale pre-trained model that can be used for various applications, typically in computer vision or natural language processing. Foundation models are usually trained in a self-supervised or unsupervised manner using approaches such as contrastive language-image pre-training (CLIP) [16], Mixup-CLIP [17] or knowledge distillation with no labels (DINO) [18, 19]. Various foundation models have been developed or applied for tasks in medical computer vision, and some recent studies have explored their use in radiological image retrieval [12, 13, 20]. However, to our knowledge, no studies have thoroughly investigated the retrieval performance of both radiological and non-radiological medical images using multiple recently developed foundation models. This represents a significant gap in research, as understanding how different foundation models perform across diverse medical imaging domains is crucial for developing robust retrieval systems. In addition to this gap, another understudied aspect of medical image retrieval, particularly for foundation models, is the effect of image size on retrieval performance. Most foundation models were originally trained on $224\times 224$ images, and the impact of using smaller images on their general performance has not been well-investigated.
@@
In this study, we conducted a comprehensive analysis of CBMIR for various types of two-dimensional (2D) and three-dimensional (3D) radiological and non-radiological medical images, using both pre-trained CNNs (as benchmarks) and recently developed foundation models that have not been thoroughly investigated in previous studies. We evaluated their retrieval performance for zero-shot retrieval without any post-processing, fine-tuning, or retraining. Furthermore, we investigated the effect of image size (ranging from very small images with only 28 pixels in one dimension to standard 224-pixel images) on retrieval performance across eight types of 2D and 3D medical images.
In summary, the main contributions and findings of this study are as follows:
  • 1.
    We compared the CBMIR performance of well-known pre-trained CNNs and recently developed medical and non-medical foundation models.
  • 2.
    We evaluated the impact of image size on the retrieval performance of both CNNs and foundation models.
  • 3.
    We performed comprehensive experiments on a variety of radiological and non-radiological medical images in both 2D and 3D formats.
  • 4.
    Our main findings indicate that for 2D datasets, foundation models achieved superior retrieval performance compared to benchmark pre-trained CNN models. However, for 3D datasets, foundation models and CNNs show comparable performance. Our experiments on image size also revealed that while competitive retrieval performance can still be achieved with small-sized images, increasing image size leads to slightly better performance, particularly for 2D datasets.
  • 5.
    To ensure reproducibility and facilitate the application of our method to new datasets, we have made our implementation publicly available.
@@ 2 Material and Methods
The general workflow of the method is shown in Figure 1, and the detailed descriptions are as follows:
Figure 1: The general workflow of the image retrieval method. CNN: Convolutional Neural Networks.
@@ 2.1 Datasets
We used a subset of eight datasets introduced in the MedMNIST V2 study [21, 22]. We selected datasets from various imaging modalities, including colon pathology, dermatoscopy, chest X-ray, fundus imaging, breast ultrasound, blood cell microscopy, abdominal CT, and electron microscopy, to investigate image retrieval performance across a wide range of medical images, both in 2D (six datasets) and 3D (two datasets). The basic statistics of these datasets, including the imaging type (2D or 3D), number of training/testing images per dataset, number of classes, and imaging modality are listed in Table 1, and example images from each dataset are shown in Figure 2. In the latest version of MedMNIST V2, 2D datasets are provided in various sizes of 28$\times$28, 64$\times$64, 128$\times$128, and 224$\times$224 pixels, and 3D datasets are provided in two sizes of 28$\times$28$\times$28 and 64$\times$64$\times$64 pixels. Further details about each sub-dataset can be found in the respective publications as referenced in Table 1. It should be noted that we evaluated the performance of the retrieval model on the test images, and the training images were used to build the database for image search. Moreover, we used the original training and testing splits of the MedMNIST V2 datasets in this study.
@@ 2.2 Pre-trained models as feature extractors
We used both well-known pre-trained CNN models (as benchmarks) and a number of foundation models as feature extractors to compare their retrieval performance on the selected sub-datasets from the MedMNIST V2. The pre-trained CNN models include VGG19 [29], ResNet50 [30], DenseNet121 [31], and EfficientNetV2M [32], while the pre-trained foundation models comprise MedCLIP [33], BioMedCLIP [34], OpenCLIP [35, 36], CONCH [37], and UNI [38]. For foundation model selection, we chose models that were not thoroughly investigated in previous studies for various types of medical image retrieval. We excluded some foundation models such as DINO V1 [18], DINO V2 [19], SAM [39], MEDSAM [40], Virchow [41], and MAE [42], as well as pre-trained ViT-based models like the original ViT [43] or Swin Transformers [44], because prior studies have demonstrated their inferior performance in medical image retrieval compared to the models selected for this study [13, 38, 12]. The basic statistics of the selected CNN and foundation models are listed in Table 2, and a brief description of each is provided in the following.
Table 2: Basic statistics of the utilized convolutional neural networks (upper part) and foundation models (lower part) in this study. Nat. = Natural, Med. = Medical.
@@ 2.2.4 EfficientNetV2M
EfficientNet is a family of CNN models initially designed to systematically and uniformly scale all dimensions of the model (depth, width, and resolution) using compound coefficients [50]. In the original study, neural architecture search [51] was employed to find a baseline model (EfficientNetB0) consisting of convolutional and mobile inverted bottleneck (MB) convolutional blocks [52], which was then scaled up to larger models for better performance (EfficientNetB1-B7). In the second version of the EfficientNet family (EfficientNetV2), a combination of Fused-MB convolutional blocks [53] (in the initial stages) and MB convolutional blocks (in the later stages) was employed and the network was optimized for improving training speed, parameter efficiency and performance. Other techniques, such as non-uniform scaling strategy, progressive learning (gradually increasing the image resolution during training), and exploiting regularizers (dropout, RandAugment, mixup), were also utilized to further boost the performance. In this study, we utilized EfficientNetV2M, which has been shown to deliver excellent performance for e.g. image classification and retrieval [32, 54] while being computationally efficient. Like other CNN-based models, we used average pooling to convert the feature maps from 8$\times$8$\times$1280 to 1280 features.
@@ 2.2.5 MedCLIP
Vision-language models such as CLIP [16] have shown excellent performance for a variety of computer vision tasks. CLIP consists of two main building blocks, a vision encoder and a text encoder that map the input images and texts to a common latent space. Using contrastive learning, the model is trained to maximize the similarity between paired images and texts. The original CLIP model was trained on 400 million general image-text pairs (named as WIT-400M) extracted from the internet. However, in the context of medical data, finding large image-text paired datasets is very challenging. In contrast to CLIP, MedCLIP is trained with three types of datasets, including medical image-text, only medical image, and only medical text datasets. To train the model, two similarity matrices were created: one referred to as the semantic similarity matrix and the other as the predicted similarity matrix. To build the semantic similarity matrix, the MetaMap [55] approach was used to extract entities from texts and images in all three types of datasets. To build the predicted similarity matrix, a text encoder (BioClinicalBERT [^1] https://huggingface.co/emilyalsentzer/Bio_ClinicalBERT) and a vision encoder (Swin Transformer [44]) were used. The utilized image encoder was pre-trained on the ImageNet dataset. Finally, to train the model, semantic matching loss was used instead of information noise-contrastive estimation (InfoNCE) loss used in the CLIP study. The model was pre-trained using two X-ray datasets, namely, MIMIC-CXR (around 600K image-text pairs) and CheXpert (around 223K images). To extract features using the pre-trained MedCLIP model, we used the vision encoder embedding, which yields 512 features for each input image.
@@ 2.2.6 BioMedCLIP
This model is an extension of the CLIP study, specifically designed for medical data. BioMedCLIP pre-training was performed using the PMC-15M dataset, which contains around 15 million image-text pairs automatically extracted from 4.4 million PubMed articles. BioMedCLIP employs a vision transformer (ViT-B/16-224) [43] as the vision encoder and PubMedBERT [56] with an extended token size of 256 as the text encoder. To extract image features from BioMedCLIP, we used the vision encoder, which yields 512 features for each input image.
@@ 2.2.8 CONCH
Contrastive learning from captions for histopathology or CONCH is a vision-language foundation model trained on over 1.17 million histological image-text pairs sourced from in-house educational notes and the PubMed central open access dataset. Similar to pre-trained CNNs, which are trained on natural images and used for various applications such as medical image analysis, we also considered using CONCH and UNI (see next section) as feature extractors for a variety of medical image types. The dataset creation to train the CONCH model involved three main steps. First, using a YOLO V5 model [59], histological images were detected. Then a GPT-style model was utilized to split the image captions [60], and finally, a CLIP model was employed to match the images with their corresponding captions. The model training was conducted using the CoCa approach [61], which includes a vision encoder (ViT-B/16), a GPT-style text encoder, and a GPT-style text decoder. Unlike CLIP, which only uses contrastive loss to maximize cosine similarity scores for pairing images and text, CoCa is also equipped with a captioning objective to generate proper text for input images. The vision encoder of the model was pre-trained on 16 million in-house images using the iBOT self-supervised approach [62], while the language model was pre-trained using more than 550,000 surgical histopathology reports from Massachusetts General Hospital and over 400,000 PubMed abstracts related to histopathology. For our study, we used the vision encoder of the CONCH model to generate 512 features for the input images.
@@ 2.2.9 UNI
General-purpose self-supervised model for computational pathology or UNI is a large-scale visual pre-trained model, originally trained on histological images. UNI was trained using the Mass-100K dataset, which contains more than 100 million H&E-stained image patches extracted from over 100,000 whole slide images. The network training was based on the DINO V2 [19] model, a self-supervised, student-teacher-based approach for pre-training. For a given image, we extracted features using the ViT-G/14 encoder of the UNI model, which yielded a feature vector of size 1024.
@@ 2.3 Distance measurement
As mentioned in Section 2.2, we used pre-trained CNNs and foundation models for feature extraction from both 2D and 3D datasets and all these models were originally trained on 2D images. For 2D datasets, we simply used the 2D training and testing images for feature extraction. For 3D datasets, however, we extracted features from each slice of a 3D train/test volume and concatenated these features to obtain the feature map for the entire 3D volume. While this approach can be computationally expensive, we selected it because our initial results indicated that applying feature reduction techniques such as principal component analysis (PCA) [63], autoencoders [64], t-Distributed Stochastic Neighbor Embedding (t-SNE) [65], or uniform manifold approximation and projection (UMAP) [66] led to a decrease in performance (refer to Table S5 and Table S6 in the supplementary materials). To measure the distance between the extracted features of the query images/volumes and the stored features in the database, we used the cosine similarity index, as suggested by previous studies on medical image retrieval [12, 16]. After calculating the index for all images (query samples and all samples in the training set), we ranked them, and the top-$k$ most similar images were selected as the output of the retrieval model.